Houghton Mifflin World Wide Web Server Proposal

net.Genesis Corporation

May 6, 1994

PROPOSAL  Houghton Mifflin

Contents

1 Introduction/Background 1 1.1 Company Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Corporation Structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Staffing and Resource Access/Availibility . . . . . . . . . . . . . . . . . . . . . . . . 2

1.3.1 Staffing and Task Allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3.2 Resource Access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 Proposal Introduction 4 2.1 General Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.2 Acknowledgement of Need . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.3 System Purpose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 Proposed System Design and Implementation Issues 6

3.1 Recommendations on Proposed System Design . . . . . . . . . . . . . . . . . . . . . 6 3.1.1 The College Division Home Page . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.1.2 Catalog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.1.3 Newsletter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.1.4 Internet Resources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.1.5 How To Use This Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

3.2 Additional Feedback Mechanisms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.3 Technology Transfer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.4 Server Set-up Implementation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

3.4.1 Hardware Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.4.2 Software Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

4 Implementation & Deliverables 13 4.1 Requirements Document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.2 Project Phases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.3 Gantt Chart . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5 Intellectual Property Rights 16

5.1 Rights to Project Information/Classified Material . . . . . . . . . . . . . . . . . . . . 16 5.2 Software Provided as Derivative License to Houghton Mifflin . . . . . . . . . . . . . 16

6 Pricing 17

6.1 net.Genesis Corporation Costs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 6.1.1 Contractor Costs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 6.1.2 Itemized Cost by Phases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 6.1.3 Training Costs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 6.1.4 Support and Maintenance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

6.2 Project Fees and Expenses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 6.3 Changes to Requirements Document during Project . . . . . . . . . . . . . . . . . . 19 6.4 Hardware And Software Costs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.5 Net connection Costs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.6 Payment Schedule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.7 Termination . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

A Itemization of Tasks 22 A.1 Server Set-up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 A.2 Software Creation/Book Catalog Phase . . . . . . . . . . . . . . . . . . . . . . . . . 22 A.3 Technical Training . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 A.4 Non-Technical Training . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 A.5 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

B Resumes 24

1 Introduction/Background

The following is an in-depth proposal for the complete development and implementation of the Houghton Mifflin College Division World Wide Web system.

1.1 Company Information

net.Genesis Corporation (net.Genesis) is a newly formed, highly motivated company devoted to effectively integrating Internet-based services into a wide range of corporate structures. net.Genesis’ primary focus lies in the World Wide Web (WWW) and NCSA Mosaic, its most prominent navigational application. We specialize in developing advanced modular systems for companies seriously interested in conducting business on the Internet.

The World Wide Web has shown itself to be the first commercially exploitable Internet-based resource. The Web’s power, manifested in its ease of use and hypermedia capabilites, can be uniquely harnessed to address the specific business needs of individual clients. Web-based systems provide an intriguing and profitable forum for advertising and sales throughout the rapidly expanding on-line community. These systems, implemented on a global scale, will augment existing markets as well as help create entirely new markets.

net.Genesis plans to be a major force in the growth and development of the Internet and the proposed National Information Infrastructure. Our goal is to develop innovative, tightly integrated WWW systems which provide corporations with an unprecedented level of market specificity, customer tracking, and system flexibility. net.Genesis is characterized by its sensitivity to the Internet culture, its responsiveness to new technology, and, most importantly, its undivided attention to their client’s business environment and business needs.

Although net.Genesis is currently holding informal discussions with several other potential clients, Houghton Mifflin would be net.Genesis’s first major client. We believe this will establish an ongoing, special relationship with Houghton Mifflin that will be greatly beneficial to both companies in the future.

In considering net.Genesis, the strength of our individual members should be noted. The net.Genesis team has the technological skills to provide world-class service, close proximity to the forefront of research in on-line areas (especially ongoing involvement in the MIT Media Lab and Laboratory for Computer Science), and a deep motivational commitment to excellence. We have exhibited our abilities in our net.Genesis demonstrations and through various WWW projects including the MIT SIPB system. net.Genesis strives to be a world-class corporation which is dedicated to serving its customers by understanding their business environment and business needs.

1.2 Corporation Structure

net.Genesis currently consists of five members: Rajat Bhargava, CEO; Matthew Cutler, President; Craig Wisneski, Director of New Media; Matthew Gray, Principal Software Architect; and George Hadjiyiannis, Principal Hardware Architect. The small number of individuals involved allows net.Genesis to be an extremely efficient, mobile, responsive, and directed corporation capable of delivering high-quality systems to exact customer specifications.

• Rajat Bhargava, CEO, is responsible for extensive business and management operations while still significantly contributing to all technical aspects of the business. He has previously done engineering work for Intel and plans to obtain a M.S. in Electrical Engineering/Computer Science (EECS) in June 1995.

• Matthew Cutler, President, provides business direction and an extensive computer background that mainly encompasses system administration, design, and application-based support. He has served as a network adminstrator for Aetna Life & Casulaty and plans to receive a B.S. in Mechanical Engineering from MIT in June 1995.

• Craig Wisneski, Director of New Media, promotes future business and market direction and is involved in technological design with a humanistic approach. He has done engineering work for Motorola and is currently doing research in the MIT Media Lab. He plans to obtain an M.S. in Media Arts & Sciences from MIT in June 1998.

• Matthew Gray, Principal Software Architect, is the director of our technological efforts in software/coding development. He set-up and currently maintains the MIT Student Information Processing Board (SIPB) WWW Server and has also done extensive work for the MIT Media Lab. He is the author of the World Wide Web Wanderer, the first program to provide accurate information on the total number of Web sites available around the world. His data and presence is widely known and respected in the Internet community. He plans to obtain a B.S. in Physics from MIT in June 1995.

• George Hadjiyiannis, Principal Hardware Architect, is the director of our technological efforts in hardware development. He has done extensive work for the MIT Laboratory for Computer Science in a variety of fields including advanced supercomputing. He holds a B.S. in EECS from MIT and plans to obtain an M.S. from MIT in EECS in January 1995.

1.3 Staffing and Resource Access/Availibility

1.3.1 Staffing and Task Allocation

For the proposed Houghton Mifflin World Wide Web server, each of the various aspects of the project will have a principal coordinator. Additionally, each member will be involved in several, if not all, other project components. George Hadijiyiannis will be responsible for hardware implementation; Matthew Gray will be responsible for software implementation; Craig Wisneski will be responsible for technology transfer including documentation and training; Matthew Cutler will be responsible for project coordination and serve as corporate liason; and Rajat Bhargava will be responsible for overall administration.

net.Genesis is actively recruiting capable, motivated, and well-positioned individuals to add to its roster. Web server operators with a demonstrated ability to develop innovative WWW systems with practical applications are currently targeted as potential net.Genesis employees. Our positioning at MIT allows us access to many of the most intelligent, enthusiastic people in this emerging field.

1.3.2 Resource Access

net.Genesis has access to many technological resources including workstations, computer graphics labs, and a variety of personal computers. In addition, our members maintain high profiles on the net in order to keep abreast of new developments, monitor general trends and start new ones.

Because of net.Genesis’s unique position in relation to the forefront of research and development of network technology, our team has access to some of the leading researchers in the field. This advantage will be translated into more cutting-edge product for our clients. net.Genesis members have close ties to leaders in sever al highly influential MIT research laboratories incluing the Media Lab the the Laboratory for Computer Science.

2 Proposal Introduction

2.1 General Background

The Internet, a world-wide system of over 30,000 computer networks, is widely acknowledged as the precursor to the National Information Superhighway. Currently, well over 25 million users, with access to over 2.2 million computers, access the Internet with unprecedented frequency. The approximate user breakdown is as follows: 55% business, 35% educational, and 10% government.

Internet users, whose overall ranks are growing by an estimated 10% per month, are increasingly turning to World Wide Web and Mosaic as the preferred means for navigation of the net. The World Wide Web, or WWW, is the first widely available electronic means for distributing multimedia documents (including images, audio, and video) over the net. The WWW includes the hypertext transfer protocol (HTTP), and the documents are written in the hypertext markup language (HTML). The Web’s growth has been explosive; WWW traffic has been increasing at a rate of 3000% per month and the number of Web servers increased six fold in the second half of 1993 alone. One of the keys to the WWW’s massive popularity has been Mosaic, a WWW navigator application updated and distributed freely by the National Center for Supercomputing Applications (NCSA). Mosaic provides a user-friendly front-end for travelling within the net, and handles hypermedia elements (large images, audio, and video) via a host of specialized external viewers which are optimized for handling hypermedia elements.

The combination of widespread availibility, power, and ease of use make the WWW ideal for commercial applications. Potential consumer uses include, but certainly is not limited to, general information distribution (such as catalogs), product demonstrations, just in time production techniques, multimedia presentations, online ordering, and virtual shopping malls. The rapidly growing market which comprises the Web is diverse, wide spread, young (average of approximately 26), and highly educated. At present, the only large corporations which distribute commercial information via WWW servers are largely technology based companies (including Apple, DEC, IBM, Lockheed, and Novell). In order for companies to avoid paying licensing fees for the WWW server software, they must provide a secondary means of electronic access to the WWW information (such as gopher or FTP).

2.2 Acknowledgement of Need

Houghton Mifflin’s College Division is ideally situated to exploit the Internet and WWW to its maximum potential. By providing a comprehensive WWW-based information system, the College Division stands to directly address the 35% share of net users involved in higher education. Additionally, Houghton-Mifflin’s publishing products are clearly portable to the user friendly, multimedia-based distribution network provided by the Web.

The benefits of such a system are numerous:

• tightly controlled, inexpensive advertising: accessible globally

• complete information on available texts: up to entire text; supplementary information in any

format

• targetting professors: able to access system 24 hrs/day; low (no) pressure browsing, instantly recieve desired textbook samples; complex searches on specific subjects; quick responses from sales representatives

• direct feedback from customers via imbedded email

• relieve pressure on sales representatives: higher efficiency; additional link to customers

• reduced costs: fewer mailings catalogs, and samples; smaller inventories

• possibility of creating on-demand custom texts

• Early participants viewed as technological leader

2.3 System Purpose

The purpose of the WWW server proposed by net.Genesis for Houghton Mifflin’s College Division is:

• To increase the effectiveness of Houghton Mifflin’s current marketing efforts by providing an additional distribution channel for brochures and product information.

• To gather valuable market information.

• To make a more direct connection between customers, sales reps, marketing managers, and sponsors.

• To increase Houghton Mifflin’s visibility and attract desirable authors through a demonstrated commitment to technology.

• To enhance Houghton Mifflin’s position among publisheres in the US college market.

3 Proposed System Design and Implementation Issues

3.1 Recommendations on Proposed System Design

net.Genesis has reveiwed the proposed Product Description as specified in the RFP and can implement this system in the time-frame outlined in Section V. When referring to the system proposed in the RFP, t he corresponding page number in RFP Appendix A appears in brackets.

3.1.1 The College Division Home Page

The College Division Home Page is the first page the usersees, and, therefore, is the most important page of the system. The design as specified in the RFP could benefit by the addition of one item: a link on the Houghton Mifflin College Division title logo that po ints to more information about the College Division. This link would include information about the Colle ge Division’s history, mission, and possibly a FAQ (listing of Frequently Asked Questions).

3.1.2 Catalog

The first page of the Catalog [2-3] has a linked index of various disciplines. The top of this page is a n excellent place to implement a keyword search. After clicking on ’search for a specific book’, the user would see a search form that can freely search the entire database. Potential searches would include any boolean combination of title, author, subject, and keyword. When a query is submitted, the user willreceive listings of the different matches to the search and be provided with links to the corresponding information.

After choosing a specific discipline, the user is presented with a more specific list of subjects withinthe desired discipline [4]. Here, the user should again have the option to conduct a search. Also, here and throughout the entire system, we propose the addition of another ’action icon’ that will take the user back one level. For example, here, the ’back’ icon would take the user from the specific discipline page [4] to the Catalog [2-3]. This is proposed for the user’s convenience, for some versions of Mosaic do not have easy access to the ’Back’ function.

After choosing a specific discipline, the user will see the resulting list of titles in that area [5]. This title list could potentially be very large. Therefore, the system shoudl have the ’action icons’ atthe top of the page. The ’back’ icon should accordingly be among these icons. Also, the option for a database search would be presented.

General Issues related to the online catalog

Entry of book information into the server Updating the contents of the catalog can be accomplished in two ways:

• Enter the majority of the necessary server information (all data pertaining to a specific text) in to the database to allow a custom converter to read the contents of the database and

automatically update the HTTP server as necessary.

• Preferred: Enter all necessary data on a new text into a custom application to be licensed from net.Genesis. The custom application automatically updates the HTTP server.

Use of Multimedia Distribution of video in the form of MPEG files is proposed but caution is advised. Quality of MPEG file s can be low depending on the contents and the method used to obtain them. Since the impression createdby low quality video is generally assumed to reflect the nature of the server, caution should be exercise d in selecting and preparing these video files. Note also that MPEG players are not real-time synchroniz ed, meaning that a 1 minute video clip on a medium performance workstation will run much slower (and be u npleasing) on an average PC and much faster on a high-end engineering workstation. Furthermore, MPEG fil es tend to be large, especially if they are of high quality. An MPEG file lasting 1 minute played at its intended speed occupy over 1.7MB of memory. Transfer of many of such files simultaneously would resultin saturation of the partial T1 outgoing line. Problems are also created when such large volumes of data are transferred over low speed lines on the client’s side. A 1.7MB file transferred over a 56Kb/s linewould take approximately 9 minutes to download. We therefore recommend that a sequence of smaller MPEG f iles is implemented, as opposed to a single large file.

Distribution of audio is more straightforward. Audio players are real-time synchronous, and while the size of high quality files is still large, they are smaller than MPEGs of corresponding time. A 1 minute au dio file would take on the order of 2.5 minutes to transfer over a 56Kb/s line.

Order Forms Currently, the secure encryption of data within Mosaic is problematic. This applies specifically to sens itive financial information, such as user credit card numbers. Adequate security requires the use of a public key encryption algorithm. The only widely available is encryption application is PGP, which uses the RSA algorythm. Unfortunately, the Federal goverment has imposed a restriction on the export of this software. Furthermore, use of PGP would be inconvenient to system users. The issue of data security is aprimary concern for the entire on-line community. A great deal of research is currently occuring in this area. The following are net.Genesis’ present proposed solutions:

• Give the option to the user to not use any protection. Thousands of users perform unprotected remote login procedures on a daily basis. During such a procedure the username and password is transmitted over the net without protection. Retrieving such information from the net is sufficiently hard that most users are not overly concerned about security risks.

• Use PGP software to encrypt the credit card information. If PGP is not available at the current site, confirm the location of the site and allow the client computer to download the software from the Houghton Miflin site. This is inconvenient and might require client PGP installation by the site administrator (as opposed to the client user).

• Submit the order form without any credit card information. When the order form is received a sales representative would call the user to confirm the order and obtain the necessary credit card information. This scenario requires a call for every submitted order form.

• Preferred method: have the user submit a request for a Houghton Mifflin ordering account before an order can be submitted. This would involve a sales representative calling the user back to obtain a credit card number, an e-mail address, and then issue a unique account number to the user. From then on, theuser inputs the account number (and not a credit card number) when placing orders. Any order made to a particular account would automatically mail the order back the the user’s predefined e-mail address. In this scenario, in the event that an account number is stolen (electronically or by other means), the original client will immediately receive notice of false orders being placed to his account.

Any of the above security strategies can be implemented by net.Genesis.

Comment Card The Comment Card will send a pre-formatted form to the Houghton Miflin server, which will in turn automatically mail it to the appropriate sales representatives as well as archive the comments.

Software Demonstrations The Software Demonstrations would run on the Houghton Mifflin server. net.Genesis can provide such functionality during the initial implemenation, but, due to server load considerations, it is recommended to defer this feature until the system has expanded to more than asingle server.

Conferencing The Conferencing area of the system would provide an interactive “chat” room (or rooms) where users could talk with textbook authors and representatives from Houghton Mifflin. net.Genesis presently possesses software for on-line conferencing that would be modified for use by Houghton Mifflin.

News Service We recommend that the News Service be based on a mailing-list system. This configuration would allow full control of access to this service; all clients would have read access to its contents, but Houghton Mifflin would dictate who updates and revises the service. An archiving system could be set up for such a service. net.Genesis can also implement other newservice structures.

3.1.3 Newsletter

The Newsletter is perhaps the most interesting part of the system for users not specifically interested in texts. The page provides useful links to high-interest information regarding the server, Houghton Mifflin College Division, and Education. The Special Bonus concept is an exciting way to attract users to the system, but we advise extreme care in selecting the shareware/freeware to be distributed through this potentially very popular item. Also, we would include another ’action icon’ (similar to those in the catalog) that will take the user back one level. Here, the icon would take the user back to the Home Page.

3.1.4 Internet Resources

The Internet Resources is a useful mirror of information contained elsewhere. We want the Houghton Miflin College Division WWW system to be viewed as a complete educational resource publically available on the Web. Through including links to as many other related systems as possible (which perhaps also provide links back), the system will become tightly integrated with existing educational information sites.

3.1.5 How To Use This Server

This page provides a simple “how to navigate the system” instructions list along with a general outline of the system layout. Important among these are the directions to disable the automatic loading of in-l ine images, which greatly affects the response time of Mosaic clients running at low (less than 56 Kb/s)speeds.

3.2 Additional Feedback Mechanisms

Houghton Mifflin will be able to obtain extensive marketing information. net.Genesis will providecustom logging software to log accesses of users to the system. The logging software will compile data on the number of accesses, location of users, and specific sites the user accessed. This software will provide key information summarizing exactly how, when, and by whom the system is accessed. This summary data will reveal general system trends, thus illustrating potential marketing directions as well as indirect suggestions for improving the system. For example, if many professors from a single school are accessing the server in a relatively short period of time, the sales representative in that region would be sent that information and respond accordingly. Additionally, if a single book (or type of book) is being heavily accessed, then the marketing and sales staffs might decide to further push that text with other means of advertising.

The logging software provided by net.Genesis is the exclusive property of net.Genesis and will remain so. Houghton Mifflin will be given a derivative license to the modifications made to the net.Genesis logging software.

3.3 Technology Transfer

net.Genesis will provide extensive documentation regarding all work completed. This will provide smooth technology transfer to Houghton Mifflin’s staff. Documentation will be provided for all key implementation phases. The documentation will be generated by the individual responsible for each task, and then reviewed by net.Genesis staff. The resulting documentation will then be transferred to the appropriate Houghton Mifflin staff for thorough revie . Documentation will be generated and reviewed throughout the project’s duration. In addition, several major documents will serve as key milestones for the project. net.Genesis acknowledges that comprhensive documentation will be vital to transferring system maintenance to the Houghton Mifflin technical staff.

Training will also be an integral aspect of the net.Genesis implementation. net.Genesis will provide extensive training in all system aspects, including both technical and non-technical staff. Training will be conducted early in the project on the Server setup and HTTP installation. The training sessions will be conducted by Matthew Gray and George Hadjiyiannis. The training session will primarily befor technical staff who will be involved in maintaining the system’s hardware.

Further training for the technical staff will be done nearer to the end of the project. The topics to be covered include the customized software that will have been written for Houghton Mifflin. If desired by the technical staff, Matthew Gray will provide an introductory seesion in PERL. Also, net.Genesis willprovide a session describing the complete system and the design choices made by the team. Again, the bulk of the technical training will be done by Matthew Gray and George Hadjiyiannis. Rajat Bhargava will also helping with this training.

Non-technical training will be held in two different sessions during the project period. Early in the project a basic class teaching HTML coding will be held. This will teach non-technical staff how to create basic WWW pages. The second training session, scheduled for week 9 of the project, will review maintaining and updating the complete working system. At this training session users will be introduced to the system and its functionality. An integral part of this training session will be the feedback from the attendees concerning ease of use, visual appeal, etc. Craig Wisneski will be project coordinator for the training of non-technical staff

net.Genesis deisres to keep the training training sessions two way; we believe that we can learn agreat deal about our implementation from the trainees as well as teach them a great deal about the subject matter. The feedback from the training sessions be critical to improving the overall system look and feel.

3.4 Server Set-up Implementation

3.4.1 Hardware Configuration

net.Genesis proposes the current implementation to consist of a single SS5/70 Sparc unit. The unitwill be equipped with a 15” color monitor (for setup and control purposes), a DB25 Ethernet interface, 64 MB DRAM (to increase performance), 535MB SCSI interface as a local disk, two external 1.05GB SCSI disksfor the server data, tape backup unit for nightly backups and a UPS (uninterrupted power supply) to protect from power brownouts. The unit will be connected through a partial or full T1 connection to the Internet. It is recommended that this connection be kept distinct from the connection that Houghton Mifflin already uses for the rest of their system for two reasons:

• Security: If the server unit is compromised it will not affect the rest of the system.

• Load isolation: The server connection is likely to become very heavily loaded on occasion. To avoid slowing net access for the whole of Houghton Mifflin’s system, it would be beneficial to keep this asa separate connection. Access to the rest of Houghton Mifflin’s network will occur by indirecting through the Internet.

A fractional T1 is recommended for the following reasons (at 384 Kb/s):

• The server’s capacity will be limited mostly by the ability to transmit data to the net. This will be especially true if heavy use of images, video and audio is made. T1 has more capacity than most other options

• Expandability: A fractional T1 is a relatively low cost alternative to a full T1. As server demand rises, more bandwidth can be purchased with minimal additional costs to support the increased demand. This expansion can be done quickly and without much effort until the demand for bandwidth exceeds the maximum for T1. At 1.5Mb/s, the full T1 bandwidth is projected to outlast the unit serving it.

Finally, the unit configuration was selected for the following reasons:

• The SS5/70 is the most cost effective high-end engineering workstation in its class. It performs only marginally lower than a SPARC 10 and costs substantially less.

• Flexibility: The combination of a UNIX type operating system and large number of peripherals makes the SS5/70 a very flexible station. Compatibility with any UNIX applications is easy to achieve and thelarge line of SUN peripherals makes it possible to adapt the system to the client’s precise needs. It also has available a number of upgrade options, such as a full expansion to a SPARC 20 machine.

• Many useful built-in features: it comes with a standard SCSI hard disk of 535MB which can easily be used to contain all the system files, software, applications, and user accounts. The SCSI interface can also be used to connect to external disks to increase the capacity to that required to hold all the catalog information.

• The extra memory is considered necessary for performance reasons. The server will be transferringlarge amounts of data to and from the hard disks. To make this process efficient, a large memory size isrequired.

• External disks: The capacity of the internal SCSI drive is insufficient to hold the vast volume of data required for the catalog. These will be stored on two external SCSI disks of a total of four timesthe capacity of the internal disk. These are projected to be of sufficient capacity for at least 2 years. Due to the small physical size of the SS5/70 it will not be possible to use internal instead of external disks.

• Tape Backup unit: the system as described above will be capable to back up the contents of its hard disks over the network to any existing tape unit. However, given the fact that the filesystem for the server will be much larger than the filesystem used for the rest of the system, it is preferable to either move that unit to the server or purchase a new one. Backups are necessary to guarantee that the system will survive!

• UPS: Power in large buildings is usually not reliable enough for vital computer systems systems. Many small power outages (on the order of a few seconds) occur regularly. A UPS will guarantee that the server will survive such outages.

3.4.2 Software Configuration

The basic platform of software over which the server will be built will be the Solaris Operating System and the Sun INS networking software. On the platform, net.Genesis will install FTP and Gopher servers and the HTTP server. More custom software will be installed to allow the services to be provided by the HTTP server. Some this software will be licensed from net.Genesis. Software in this category includes, but is not limited to, the conferencing software, the HTML editors, the HTML converters, auto-reply systems, and logging software. The remaining software will be tailored on site to interface betweenthe licensed software and the existing Houghton Mifflin systems. All software will be designed in independent modules facilitating easy expansion and modification should the need arise. On-line documentationwill be provided for most of the software and will be incorporated into existing on-line documentation.

4 Implementation & Deliverables

An itemized list of tasks, divided by phases, has been included as Appendix A.

4.1 Requirements Document

Upon initiation of the project, one of the first tasks net.Genesis will undertake is the development of a requirements document. The document, to be co-produced with key Houghton Mifflin participants in the project, will specify the exact definition of the system to be developed by net.Genesis. The definition will specify deliverables and milestones for the project. The requirements document will be signed by key executives from both parties. Upon completion of the document net.Genesis reserves the right to reevaluate the price structure should the definition of the project change from the RFP/Proposal until then. However, only a timetable change of greater than 10% will warrant a material change in the price.

4.2 Project Phases

Server Setup Phase

A workstation will be obtained and configured on site. All necessary software will be installed, including NFS, the web server, and necessary UNIX utilities. A file system for the HTTP server data (mostly the catalog information) will be created. At the same time, accounts for the Houghton Mifflin technical staff and the net.Genesis staff will be created and access to the unit over the network will be provided (for testing purposes). An NFS mount point connecting the unit’s file system to that of the rest of the Houghton Mifflin system (including the database) will be established.

Deliverables

• Server configured

• Web server and other software installed

• Accounts Created

• NFS mount point established

Software Creation and Book Catalog Phase

We will evaluate Houghton Mifflin’s existing database and will implement necessary software to convert the available online information into hypertext format for use on the Web server. A hypertext document heierarchy will be established for the book catalog and implemented in the conversion software. A search engine interface to the book catalog will be implemented including an interface via the Web server. Any further documents and pages to be presented on the Web server will be converted to HTML. Additionally, log analysis software will be provided to examine server usage patterns and growth and conferencing software will be provided to allow MUD-like interaction on the system. Full documentation on all conversion software will be provided.

Deliverables

• Software package for conversion from native DB format to HTML documents

• Software package for easy modularization of hypertext documents

• Licencse for log anylsis and conferencing software

• Search engine software

• Complete conversion to HTML of all necessary documents for the Web server

• Complete documentation

Technical Training Phase

Full training of technical staff will provided in all areas required for Web server maintenance, including server internal software, HTML and HTTP specifications, and all conversion software being provided. Specific technical training will be provided for the licensed software including modifications. Also, training in the techniques and programming languages used to create the custom software (software developed on site) can be included. Reasonable technical knowledge and familiarity with basic programming techniques and languages will be assumed.

Deliverables

• Technical Staff Training Seminars

• Technical Staff documentation

Non-Technical Training Phase

Non-technical staff will be familiarized with usage of the web server, performing updates on existing documents, and in the basics of writing hypertext documents. Use of application programs for catalog update, conferencing software, and other related systems will be covered. Also, the basics of HTML can be overviewed.

Deliverables

• Non Technical Staff Training Seminars

• Non Technical Documentation Support Phase net.Genesis will be providing support after the completion of the project. After product release, pointers will be established to the server from a variety of locations such as USENET groups, other HTTP sites, marketing information and other publicly accessible sites. This will make the server widely known. net.Genesis will also provide maintenance service which will include, but not be limited to, help on modifications, expansion, repairs and upgrades. Such service is included in the current contract for a period of 3 months up to a maximum of 200 hours. Deliverables

• Publication of Server Location

• Assistance on Modifications, Expansion, Upgrades and Repairs

4.3 Gantt Chart

Key Milestones as listed are:

• Requirements Document

• Fully Configured Server (including HTTP installation)

• Training of Technical Staff in Server Internal Software and HTTP

• Specifications Document for Book Catalog System

• Fully Functional HTML Converters

• Fully Functional Search Engine

• Customized Log Analysis and Conferencing Software

• Operational Server with Book Catalog and Features

• Completion of Feedback and Updating of Server and Catalog

• Operational FTP/Gopher Mirror Sites

• Final Release of WWW Server

5 Intellectual Property Rights

5.1 Rights to Project Information/Classified Material

All reports, documents, instruments, data, or material which are received by net.Genesis from Houghton Mifflin or developed by net.Genesis in connection to the services provided under this agreement, except those under derivative license, are, and will continue to be, the property of Houghton Mifflin and will not be divulged, distributed or released to any third party without the consent of Houghton Mifflin except as may be required by order of any Court of competent jurisdiction. All software material and documentation created by net.Genesis prior to work in conjunction with Houghton Mifflin will continue to be the exclusive rights of the creator, net.Genesis. Houghton Mifflin shall be granted a derivative license for modifications done at Houghton Mifflin while using software code priorly developed by net.Genesis. The derivative license will allow Houghton Mifflin the use and modification of the derivative software. Under no circumstances ! may Houghton Mifflin divulge, dist ribute, or release the derivative work to any third party without the express written consent of net.Genesis.

5.2 Software Provided as Derivative License to Houghton Mifflin

• Log Analysis Software

• Conferencing Software

• Proprietary Search Engines

• Automated Reply

• Custom HTML Converters/Editors

• WWW Server Utility Routines

• Order Form Processing

6 Pricing

6.1 net.Genesis Corporation Costs

6.1.1 Contractor Costs

Based upon the complexity and time required to implement the system proposed by net.Genesis to Houghton Mifflin, the final cost has been determined to be $129,600. net.Genesis recognizes the factthat Houghton Mifflin is our first customer and would like the price to reflect this fact. We have provided Houghton Mifflin with a 20% discount off of our projected rate of $162,000. Furthermore, the payment schedule has been defined to reward net.Genesis upon completion of key milestones. This helps to minimize Houghton Mifflin’s financial risks and also gives net.Genesis the incentive to provide a world class product on schedule. The total cost for the system as proposed in this document, subject to review by the Requirements document, will be $129,600. The above price is subject to change pending outcome of the Requirements document. Changes affecting the timetable in a material way will render the above figure void.

Total Cost of net.Genesis Contract Work $162,000
Less Houghton Mifflin Discount (20%)  ($32,400)
Total Cost  $129,600

Costs are based upon 50 man-hour work weeks for the life of the project (10 weeks). 200 man-hours of support of the system over a 3 month period beginning upon completion of the total project is included in the total cost the system over a 3 month period of time is included in the total cost. If support beyond 200 man hours is desired additional charges of $60/man-hour will be charged. Further support options are outlined in the Support section below.

Total Costs Include:

• Comprehensive review of requirements and specifications

• Server setup and integration

• HTTP installation

• FTP/Gopher installation

• Lifetime licence for net.Genesis proprietary software

– HTML Editors/Converters

– Conferencing software

– Log Analysis software

– Order form processing

• Creation of custom World Wide Web pages • Integration of audio/video/graphic design into WWW pages

• Training

• Complete documentation of system

• Technology transfer

• Support and maintenance

6.1.2 Itemized Cost by Phases

Initial Phase  $13,900

Requirements Document Server Setup and Test Server Integration with LAN HTTP Installation Full Documentation

Book Catalog Phase 1  $28,900

Evaluation/Planning/Design of System Specification/Creation/Testing of Custom HTML Converters Full Documentation

Book Catalog Phase 2  $33,800

Spec/Creation/Test of Search Engine/Conferencing/Log Analysis Creation of Pages Conversion and Quality Control of Data Full Documentation

Book Catalog Phase 3  $27,520

Brining server on-line Feedback and updating of Server/Pages FTP/Gopher Mirror Sites Final Release of Server Full Documentation

Training  $15,880

Technical Training Non-Technical Training

Support and Maintanence  $9,600

200 man-hours of support and maintenance for 3 months 6.1.3 Training Costs

On site training costs for the duration of the project have been included in the total cost. Training during the project will include Houghton Mifflin’s technical and non-technical staff. Any off-site training is subject to additional contractor charges. Expenses such as air travel, hotel, rental car, etc. will be paid by Houghton Mifflin if off-site training is requested.

6.1.4 Support and Maintenance

After the 3 month support period has expired, Houghton Mifflin can obtain a service contract for the duration of 1 year for $17,500. The service contract will entitle Houghton Mifflin to 20 man-hours of support per month. Additional service can be purchased on an hourly basis. Phone/email correspandance willbe charged fees of $70/man-hour. On site support will be charged at $100/man-hour. net.Genesis fully understands that system downtime translates directly into compromised business oppurtunities, whether those be actual lost orders or loss of customer satisfaction. Accordingly, net.Genesis pledges to provide complete, prompt service whenever it should be necessary.

6.2 Project Fees and Expenses

Project Costs above include only direct contractor fees. Not included in the fees are any expenses net.Genesis incurs during the course of the project related to the project. These may include, but arenot limited to, workstation rental, outside contract work regarding conversion of data, the procurement of any necessary documentation, and phone, mail, fax, copying exclusively related to the Houghton Mifflinproject. Houghton Mifflin will be responsible for reimbursing net.Genesis for all direct project e xpenses incurred to net.Genesis.

6.3 Changes to Requirements Document during Project

Changes made after the requirements document has been signed by both parties will be subject to additional charges.

6.4 Hardware And Software Costs

Hardware Costs SS5/70 SPARC workstation (15” color monitor, CG3 Video Frame Buffer,

16MB DRAM, 535MB SCSI Hard Disk)  $3,995
Additional 48MB DRAM  $3,450
2x1.05GB SCSI External Expansion Disks  $2,500
5.0GB 4mm Tape Desktop Backup Unit  $1,600
Uninterruptible Power Supply  $300

Software Costs

Solaris 2.3 Multi-user license  $300
C compiler  free
FTP/GOPHER  free
net.Genesis software (derivative license)  included in contract

6.5 Net connection Costs

A full or partial T1 connection is recommended with the vendor providing nameservice:

Bandwidth  Annual Fee Com. Routing Installation
128 Kb/s  $12,550  $3,000  $14,950
384 Kb/s  $15,850  $3,500  $14,950
768 Kb/s  $20,800  $4,500  $14,950
T1(1.54 Mb/s) $27,400  $6,500  $14,950

Installation costs may vary depending on how much of the installation the client is willing to do for themselves.

6.6 Payment Schedule

Payments to net.Genesis shall be payable in installments according to key milestones in the implementation process. Final dates will be determined upon completion of the Requirements Document, which will be signed by key executives of both companies.

Project Launch  20% $25,920
Milestone 1  10% $12,960

Requirements Document Server Setup Training for Server/HTTP

Milestone 2  10% $12,960

Creation of periphery pages Creation of custom editors (tested)

Milestone 3  30% $38,880

Completion of all pages Completion of search engine, conference feature, logging software, usenet group access Bring WWW Server up and on line (ready for internal introduction)

Milestone 4  30% $38,880

FTP/Gopher Mirror Sites Final Release (to net)

Total  $129,600

Terms will be as follows. Invoices will be delivered to Houghton Mifflin upon the completion of the milestones outlined in this document and specified in the requirements document. Houghton Mifflin will have30 days to submit payment. A finance charge to 1.5% per month will be applied to all overdue invoices.

6.7 Termination

The contract may be terminated by Houghton Mifflin, with or without cause, immediately upon written notice and the payment of a termination fee of 20% of total project fees or the remaining amount of the total project fee, whichever is the lesser. The remaining project fee is equal to the total agreed project fee less outstanding and paid invoices. Expenses will be paid as incurred separate from project fees.

A Itemization of Tasks

A.1 Server Set-up

• Obtain workstation and internet connection

• Determine logistical questions about server

• Determine optimal integration scheme with Houghton Mifflin LAN

• Determine adequate level of security for different phases of implentation

• Configure of OS

• Configure of file system

• Create unix accounts

• Install of NFS software (and integration with current LAN)

• Install of logical partitions and logical disks

• Install internet access

• Install LAN access

• Install HTTP

• Install other binaries

• Testing of server (security)

• Full Documentation

A.2 Software Creation/Book Catalog Phase

• Evaluate Houghton Mifflin’s current database

• Evaluate file formats

• Evaluate state of audio/video

• Find high quality quicktime to MPEG converter

• Evaluate graphic design needs

• Evaluate Lotus Notes email system for Replies

• Determine security features necessary

• Determine file structure • Determine search engine features

• Determine process of conversion of audio/video files

• Quality control process for audio/video

• Determine necessary editors/converters and write specification

• Create custom HTML converters

• Convert audio/video files

• Quality control

• Create GIFs

• Create icons from graphic design work

• Create actual book pages

• Create periphery pages

• Create main menu pages

• Create search engine

• Create conference feature

• Create usenet functionality

• Create logging programs

• Test all converters

• Exhaustive tests to ensure correct conversion

• Test pages, search engine

• Test from remote sites

• Full documentation

• Feedback from Houghton Mifflin

A.3 Technical Training

• Documentation management

• Creation of training materials for mini-courses

• Training sessions

• HTML

net.Genesis Corporation • HTTP

• Server setup and mantainance

• Script/Converter

A.4 Non-Technical Training

• Development of Training Materials for Mini Courses

• HTML

• Updating System

• Use of System

A.5 Support

• Hel Houghton Mifflin technical and non-technical staff with questions and problems regarding work completed by net.Genesis

• Product release – Internet announcements

– Usenet announcements

– Incorporation into Marketing Information

B Resumes